Mathematics 244 – Linear Algebra Definitions Quiz 5 – March 2, 2007

Your Name:

A) (5) Let A be an n × n matrix. Give the expansion of the determinant of A along the first row, and explain exactly what the notation A_{1j} means.

B) (5) True or False, and explain (briefly!): If A is a 5 × 5 matrix with det(A) = 0, then the system Ax = b has infinitely many solutions for all b ∈ R^{5}.